OpenStudy (anonymous):

What is the equation in point-slope form of the line passing through (−2, 1) and (4, 13)? y − 1 = −2(x − 4) y − 1 = 2(x + 2) y − 13 = −2(x + 2) y − 13 = 2(x + 4)

OpenStudy (anonymous):

y − 1 = 2(x + 2)

OpenStudy (anonymous):

by two point form as if we have point let (a,b) and (c,d) then equation of line is (y-b)=((b-d)/(a-c))(x-a) so here (-2,1) and (4,13) y-1=((1-13)/(-2-4))(x-(-2)) => y-1=2(x+2)